### v2

# A note on internality of certain differential systems

Abstract.
We prove two results, generalizing certain theorems by Jin and Moosa [JM20], on the internality of the system of differential equations
[…]
where $f$ and $g$ are rational functions in one variable.
Let $K$ be a differential field of characteristic zero. We fix a universal differential field extension [^1] A differential field extension $U$ of $K$ is said to be an universal extension of $K$ if for any finitely generated differential field extension of $F$ of $K$ contained in $U$ and for any finitely generated differential field extension $F_{1}$ of $F,$ which is not necessarily contained in $U$, there exists a differential $F$-embeddeding from $F_{1}$ into $U$ (see [Kol73, page 133]). $U$ of $K$ and assume that all differential field extensions of $K$ considered in this article are contained in $U.$ For any differential field extension $M$ of $K,$ the subfield of constants of $M$ is denoted by $C_{M}.$ A generic solution over $K$ of a system of first order differential equations
[…]
where $f_{1},\dots,f_{n}$ are rational functions in $n$ variables over $K,$ is a tuple $(y_{1},\dots,y_{n})\in U^{n}$ such that for each $i,1\leq i\leq n,$ $y^{\prime}_{i}=f_{i}(y_{1},\dots,y_{n})$ and the field transcendence degree $\mathrm{tr.deg}(K(y_{1},\dots,y_{n})|K)=n.$
[…]
The model theoretic concept of internality of a system of differential equations is known to be closely related to Kolchin’s strongly normal extensions (see [Poi01, Section 2.7]), and it has its origin in the works of Rosenlicht [Ros74, Proposition] and Zilber [Zil77]. In [JM20], Jin and Moosa prove two theorems, which appear as Theorem A and Theorem B in their paper, on almost internality of the system of differential equations
[…]
where $f$ is a rational function in one variable. In Theorems 4 and 5 of this article, we generalize and extend both these results of Jin and Moosa. In particular, Theorem 4 completely answers the question of when the system (L) is almost internal to the constants, given that $x^{\prime}=f(x)$ is internal to the constants.
Proposition 1.
Let
[…]
be a system almost internal to the constants. Then, there exists a finitely generated differential field extension $M$ of $K$ such that for any generic solution $(x,y)$ over $M$ of the system [^2] Generic solutions over $M$ of the system always exist. To see this, consider the rational function field $M(s,t)$ and equip this field with the derivation induced by the system: $s^{\prime}=f(s,t)$ and $t^{\prime}=g(s,t)$. Then, by definition, there is a differential $M$-embedding of $M(s,t)$ into $U.$ The image of $s,t$ under this embedding provides a generic solution over $M$ of the system., the differential field $M(x,y)$ contains two $M$-algebraically independent constants. Furthermore, if $N$ is a differential field intermediate to $M(x,y)$ and $M$ with $\mathrm{tr.deg}(N|M)=1,$ then $\mathrm{tr.deg}(C_{N}|C_{M})=1.$
Proof.
Since the system is almost internal to the constants, there are generic solutions $(x_{1},y_{1}),\dots,(x_{m},y_{m})$ over $K$ of the system such that for any generic solution $(x,y)$ of the system over $M:=K(x_{1},y_{1},\dots,x_{m},y_{m}),$ there are constants $c_{1},\dots,c_{l}$ such that $x$ and $y$ are algebraic over $M(c_{1},\dots,c_{l}).$ We now show that $M(x,y)$ contains two $M$-algebraically independent constants. Let $c_{1},\dots,c_{r}$ be an $M$-transcendence basis of constants for $M(c_{1},\dots,c_{l}).$ Since $\mathrm{tr.deg}(M(c_{1},\dots,c_{l},x,y)|M)=r,$ we must have $\mathrm{tr.deg}(M(c_{1},\dots,c_{l},x,y)|M(x,y))=r-2.$ Therefore, there is a nonzero polynomial $P\in M(x,y)[X_{1},\dots,X_{r}]$ such that $P(c_{1},\dots,c_{r})=0.$ Since $c_{1},\dots,c_{r}$ are constants, it follows that such a polynomial $P$ must belong to $C_{M(x,y)}[X_{1},\dots,X_{r}]$ ([Kol48, Section 14, Theorem 2]). As $\mathrm{tr.deg}(M(c_{1},\dots,c_{l})|M)=r,$ it also follows that one of the coefficients of $P$ must be a constant nonalgebraic over $M.$ If we call this constant $u_{1}$ and replace $M$ by $M(u_{1}),$ then a similar argument shows the existence of another constant $u_{2}\in M(x,y),$ which is nonalgebraic over $M(u_{1}).$ Observe that either $u_{1}$ or $u_{2}$ must belong to $M(x,y)\setminus M(x).$
Let $N$ be a differential field intermediate to $M(x,y)$ and $M$. Since $\mathrm{tr.deg}(N|M)=1,$ there is a nonzero polynomial $P\in C_{N}[X_{1},X_{2}]$ such that $P(u_{1},u_{2})=0.$ Again, by the same argument as above, there is a constant $v\in N$ which is not algebraic over $M.$
Theorem 2.
Let $f$ be a rational function in one variable over $K.$ The equation $x^{\prime}=f(x)$ is internal to the constants if and only if $f(x)=a_{2}x^{2}+a_{1}x+a_{0},$ where $a_{0},a_{1},a_{2}\in K.$
Proof.
Let $x^{\prime}=f(x)$ be internal to the constants. Then, there are generic solutions $x_{1},\dots,x_{m}$ over $K$ such that for any generic solution $x$ over $M:=K(x_{1},\dots,x_{m})$ there are constants $c_{1},\dots,c_{l}$ such that $x\in M(c_{1},\dots,c_{l})$. From [Kol73, Chapter 2, Corollary 2], we know that $M(x)$ is generated over $M$ as a field by a set of constants. Therefore, we may assume that $M(x)=M(c_{1},\dots,c_{l}).$ We first claim that there exists an algebraic extension $\tilde{M}$ of $M$ such that $\tilde{M}(x)=\tilde{M}(c)$ for some constant $c$.
Note that $C_{M(x)}$ and $M$ are linearly disjoint over $C_{M}$ ([Kol73, Chapter 2, Corollary 1]). Therefore, $C_{M}(c_{1},\dots,c_{l})$ and $M$ are linearly disjoint over $C_{M}.$ Since $M(c_{1},\dots,c_{l})=M(x),$ we have $\mathrm{tr.deg}(M(c_{1},\dots,c_{l})|M)=1$ and from linear disjointness, we obtain $\mathrm{tr.deg}(C_{M}(c_{1},\dots,c_{l})|C_{M})=1.$ Now since $M(c_{1},\dots,c_{l})=M(x)$ is a rational function field, there is a finite algebraic extension $\tilde{C}_{M}$ of $C_{M}$ such that $\tilde{C}_{M}(c_{1},\dots,c_{l})=\tilde{C}_{M}(c)$ ([Che63, Theorem 5]). Any extension of a zero derivation on a field to its algebraic closure is again a zero derivation. Therefore, $c$ must also be a constant. Let $\tilde{M}:=M\tilde{C}_{M}$ and observe that
$\tilde{M}(x)=\tilde{M}(c_{1},\dots,c_{l})=\tilde{M}(c).$
This proves the claim.
Since $\tilde{M}(x)=\tilde{M}(c),$ there exists $\alpha,\beta,\gamma,\delta\in\tilde{M}$ with $\alpha\delta-\beta\gamma\neq 0$ such that
$c=\frac{\alpha x+\beta}{\gamma x+\delta}.$
A simple calculation shows that $f(x)=x^{\prime}=a_{2}x^{2}+a_{1}x+a_{0}$ for some $a_{0},a_{1},a_{2}\in\tilde{M}$. Now since $f\in K(x),$ we obtain that $a_{0},a_{1},a_{2}\in K.$
To prove the converse, let $x,x_{1},x_{2},x_{3}$ be distinct generic solutions of $x^{\prime}=a_{2}x^{2}+a_{1}x+a_{0}$ over $K.$ It is observed in [Nag20, Page 102] that
$c:=\frac{(x-x_{2})(x_{3}-x_{1})}{(x-x_{1})(x_{3}-x_{2})},$
is a constant [^3] It is easily seen that the first order differential equation $y^{\prime}=(a_{1}(x+x_{1}+x_{2}+x_{3})+a_{0})y$ has $(x-x_{2})(x_{3}-x_{1})$ and $(x-x_{1})(x_{3}-x_{2})$ as its solutions and therefore $c,$ being the ratio of these two solutions, must be a constant.. Since
$x=\frac{x_{2}(x_{3}-x_{1})+cx_{1}(x_{2}-x_{3})}{(x_{3}-x_{1})+c(x_{2}-x_{3})},$
we now obtain $x\in K(x_{1},x_{2},x_{3},c).$ This proves that the equation $x^{\prime}=a_{2}x^{2}+a_{1}x+a_{0}$ is internal to the constants. ∎
Remark 3.
We note here that a particular case of Theorem 2 can be derived from certain recent results ([JM25, KRS24]). Suppose that $K$ is an algebraically closed field and that the equation $x^{\prime}=f(x)$ is both internal to the constants and weakly orthogonal to the constants [^4] That is, the rational function field $K(x)$ with the extended derivation $x^{\prime}=f(x)$ has the same field of constants as $K.$. Then, it is known that the binding group of the equation $x^{\prime}=f(x)$ is a linear algebraic group. Therefore, one can apply [JM25, Proposition 6.5] and conclude that $K(x)=K(y),$ where $y^{\prime}=ay^{2}+by+c$ for some $a,b,c\in K.$ Writing
$x=\frac{\alpha y+\beta}{\gamma y+\delta},$
where $\alpha,\beta,\gamma,\delta\in K$ with $\alpha\delta-\beta\gamma\neq 0,$ and taking derivatives, we obtain the conclusion of Theorem 2 that $f$ is a polynomial over $K$ of degree at most $2.$
The conditions that $x^{\prime}=f(x)$ is both internal to the constants and weakly orthogonal to the constants imply that the differential field $K(x)$ can be embedded in a strongly normal extension of $K.$ Now, applying [KRS24, Theorem 1.1 (iii)], we obtain that $f$ is a polynomial over $K$ of degree at most $2.$ Thus, Theorem 2 provides a generalization of these results, as it requires neither that $x^{\prime}=f(x)$ be weakly orthogonal to the constants nor that $K$ be an algebraically closed field.
Under the assumption that $x^{\prime}=f(x)$ is internal to the constants with the binding group not of dimension $3,$ it is shown in [JM20, Theorem A] that the system (L) is almost internal to the constants if and only if it splits; which implies, the system can be transformed into a system of the form
$\begin{cases}\begin{aligned} x^{\prime}&=f(x)\\ y^{\prime}&=ay\\ \end{aligned}\end{cases},$
where $a\in K.$ In the next theorem, using Theorem 2 and Proposition 1 and without any assumption on the binding group of $x^{\prime}=f(x)$, we shall completely determine when the system (L) is almost internal to the constants.
Theorem 4.
Let $K$ be a differential field and $f$ be a nonzero rational function in one variable over $K$. The system (L) is almost internal to the constants and $x^{\prime}=f(x)$ is internal to the constants if and only if $f(x)=a_{2}x^{2}+a_{1}x+a_{0},$ where $a_{0},a_{1}\in K$ and $a_{2}$ is a nonzero rational number.
Proof.
We know from Theorem 2 that $f(x)=a_{2}x^{2}+a_{1}x+a_{0},$ where $a_{0},a_{1},a_{2}\in K.$ Suppose that the system (L) is almost internal to the constants. By Proposition 1, there exists a differential field extension $M$ of $K$ such that for any generic solution $(x,y)$ over $M,$ the differential field $M(x,y)$ contains two $M$-algebraically independent constants. Then $M(x,y)$ contains a constant which is not in $M(x)$ and therefore the latter must contain an element $z$ such that $z^{\prime}=nxz$ for some positive integer $n$ ([Mag94, Example 1.11]). Let $z=P/Q$ where $P,Q\in M[x].$ We can assume that $P$ and $Q$ are relatively prime and $Q$ is monic. Write $P=\alpha_{p}x^{p}+\alpha_{p-1}x^{p-1}+\cdots,Q=x^{q}+b_{q-1}x^{q-1}+\cdots\in M[x]$ with $\alpha_{p}\neq 0.$ Then
$P^{\prime}Q-Q^{\prime}P=nxPQ$
and we obtain $(p-q)a_{2}=n.$ Since $n\neq 0,$ $a_{2}=n/(p-q)$ is a nonzero rational number.
For the converse, suppose that $f(x)=a_{2}x^{2}+a_{1}x+a_{0},$ where $a_{0},a_{1}\in K$ and $a_{2}=m_{1}/m_{2},$ $(m_{1},m_{2})\in\mathbb{Z}^{2}$ a nonzero rational number. Let $(x,y)$ be a generic solution of the system (L). We shall now find generic solutions $(x_{1},y_{1})$ and $(x_{2},y_{2})$ of the system and constants $c_{1},c_{2}$ so that $x,y$ are algebraic over the field $K(y_{1},x_{1},y_{2},x_{2},c_{1},c_{2}).$ This will then complete the proof of the theorem.
Let $v$ be an element algebraic over the field $K(x,y)$ satisfying $v^{m_{2}}=y^{-m_{1}}.$ Then $v^{\prime}=-a_{2}xv,$ $(a_{2}x)^{\prime}=a_{2}x^{\prime}=(a_{2}x)^{2}+a_{1}(a_{2}x)+a_{2}a_{0}$ and $v^{\prime\prime}=a_{1}v^{\prime}-a_{2}a_{0}v.$ Consider the rational function field $K(v,v^{\prime})(v_{1},v^{\prime}_{1},v_{2},v^{\prime}_{2})$ and extend the derivation on $K(v,v^{\prime})$ to $K(v,v^{\prime})(v_{1},v_{2},v^{\prime}_{1},v^{\prime}_{2})$ by declaring $v^{\prime\prime}_{1}=a_{1}v^{\prime}_{1}-a_{2}a_{0}v_{1}$ and $v^{\prime\prime}_{2}=a_{1}v^{\prime}_{2}-a_{2}a_{0}v_{2}$. Let $N=K(v_{1},v^{\prime}_{1},v_{2},v^{\prime}_{2}).$ Note that $v_{1},v_{2}$ are $C_{U}$-linearly independent and that the Wronskian of $v,v_{1},v_{2}$ is zero. Therefore $v=c_{1}v_{1}+c_{2}v_{2},$ for some constants $c_{1},c_{2}.$ This implies, $v\in N(c_{1},c_{2})$ and since $a_{2}x=-v^{\prime}/v\in N(c_{1},c_{2}),$ we obtain that $x\in N(c_{1},c_{2}).$
Let $y_{1}$ and $y_{2}$ be elements algebraic over the field $N$ such that $v^{m_{2}}_{1}=y^{-m_{1}}_{1}$ and $v^{m_{2}}_{2}=y^{-m_{1}}_{2}.$ Let $x_{1}:=-v^{\prime}_{1}/a_{2}v_{1}$ and $x_{2}:=-v^{\prime}_{2}/a_{2}v_{2}.$ Then, it is easy to see that $(x_{1},y_{1})$ and $(x_{2},y_{2})$ are generic solutions of the system (L). Now we consider the field $K(y_{1},x_{1},y_{2},x_{2},c_{1},c_{2}).$ Clearly, $v_{1}$ and $v_{2}$ are algebraic over $K(y_{1},x_{1},y_{2},x_{2},c_{1},c_{2})$ and since $v=c_{1}v_{1}+c_{2}v_{2},$ we obtain that $v$ is also algebraic over $K(y_{1},x_{1},y_{2},x_{2},c_{1},c_{2}).$ Since $x=-v^{\prime}/a_{2}v,$ we obtain that $x$ is algebraic over $K(y_{1},x_{1},y_{2},x_{2},c_{1},c_{2}).$ Finally, $v^{m_{2}}=y^{-m_{1}}$ implies that $y$ is algebraic over $K(y_{1},x_{1},y_{2},x_{2},c_{1},c_{2}).$
Theorem 5.
Let $C$ be an algebraically closed differential field with zero derivation, and $f,g$ be two nonzero rational functions in one variable over $C.$ The system
[…]
is almost internal to the constants if and only if the following two conditions hold:
  1. (i)
    $\frac{1}{f(x)}=\frac{\partial u}{\partial x}$ or $\frac{1}{f(x)}=\frac{c}{u}\frac{\partial u}{\partial x}$ for some $c\in C\setminus 0$ and $u\in C(x).$
  2. (ii)
    $\frac{b+mg(x)}{f(x)}=\frac{1}{v}\frac{\partial v}{\partial x}$ for some integer $m\in\mathbb{Z},$ $b\in C$ and $v\in C(x).$
Proof.
Let $(x,y)$ be a generic solution of the system (A) and $F=C(x,y).$ We first make the following observations. Let $z\in F\setminus C$ such that $z^{\prime}=0.$ Then $z\notin C(x);$ otherwise $f(x)\frac{\partial z}{\partial x}=0,$ which is a contradiction. Therefore, $z\in F\setminus C(x).$ As noted in the proof of Theorem 4, there exists a positive integer $m$ and an element $v\in C(x)$ such that $v^{\prime}=mg(x)v.$ Therefore, $f(x)\frac{\partial v}{\partial x}=mg(x)v$ and we obtain
$\frac{mg(x)}{f(x)}=\frac{1}{v}\frac{\partial v}{\partial x}.$
Thus, if $z\in F\setminus C$ such that $z^{\prime}=0$ then $z\in F\setminus C(x)$, and in that case the condition (ii) holds with $b=0$.
Along with these observations, we suppose that the system (A) is almost internal to the constants. Then, by Proposition 1, there exists a differential field extension $M$ of $C$ such that $M(x,y)$ contains two $M$-algebraically independent constants. Furthermore, since $M(x)$ is a differential subfield of $M(x,y)$ with $\mathrm{tr.deg}(M(x)|M)=1,$ it follows that $\mathrm{tr.deg}(C_{M(x)}|C_{M})=1.$ By [Ros74, Proposition], $C(x)$ contains an element $u$ such that either $u^{\prime}=1$ or $u^{\prime}=cu$ for some nonzero $c\in C.$ Then $\frac{1}{f(x)}$ is either $\frac{\partial u}{\partial x}$ or $\frac{1}{f(x)}=\frac{c}{u}\frac{\partial u}{\partial x}.$ This implies that condition (i) holds.
To prove that condition (ii) also holds, we consider the following two cases. Suppose that $C_{F}\neq C$. Then, by the observation made in the first paragraph, the condition (ii) holds. Now suppose that $C_{F}=C.$ Since there exists a differential field $M$ such that $M$ and $F$ are free over $C$ and $MF$ contains two $M$-algebraically independent constants, by [KS25, Proposition 4.2], there is a differential field $L$ intermediate to $C$ and $F$ such that $\mathrm{tr.deg}(L|C)=2$ and $L$ can be embedded in a strongly normal extension of $C.$ Since $L$ is contained in the purely transcendental extension $C(x,y),$ by [KS25, Theorem 1.2], $L$ can be embedded in a Picard-Vessiot extension of CC [^5] The existence of the field $L$ can also be obtained using results from model theory (see [EJ24, Theorem 3.5]). The differential Galois group of any Picard-Vessiot extension of $C$ is a connected commutative linear algebraic group over $C$. Therefore, by the fundamental theorem of differential Galois theory, $L$ must also be a Picard-Vessiot extension of $C$ with a differential Galois group isomorphic to $\mathbb{G}_{a}\times\mathbb{G}_{m}$ or $\mathbb{G}_{m}\times\mathbb{G}_{m}.$ This implies, $L=C(\alpha,\beta)$ where $\alpha,\beta$ are $C$-algebraically independent, $\beta^{\prime}=c_{2}\beta$ for some nonzero $c_{2}\in C$ and either $\alpha^{\prime}=1$ or $\alpha^{\prime}=c_{1}\alpha$ for some nonzero $c_{1}\in C.$ Consequently, there exists an element $\zeta\in C(x,y)$ such that $\zeta^{\prime}=a\zeta$ for some nonzero $a\in C.$ By the Kolchin-Ostrowski theorem, there exists integers $m,n$ such that $h:=y^{m}\zeta^{n}\in C(x).$ Then
$h^{\prime}=f(x)\frac{\partial h}{\partial x}=my^{m}\zeta^{n}g(x)+nay^{m}\zeta^{n}.$
From the above equation, we obtain
$\frac{mg(x)+b}{f(x)}=\frac{1}{h(x)}\frac{\partial h}{\partial x},\quad na=:b\in C.$
Thus, we have shown that the condition (ii) holds in both cases.
To prove the converse, let $(x,y)$ be a generic solution of the system (A). From conditions (i) and (ii), we obtain that $C(x)$ contains an element $s$ such that $s^{\prime}=1$ or $s^{\prime}=cs$ for some nonzero $c\in C$ and that there is an element $t\in C(x,y)\setminus C(x)$ such that $t^{\prime}=bt$ for some $d\in C.$ Note that $s=u$ and $t=y^{m}/v.$ Now let $(x_{1},y_{1})$ be another generic solution of the system (A). Consider the differential $C$-isomorphism $\phi:C(x,y)\rightarrow C(x_{1},y_{1}),$ where $\phi(x)=x_{1}$ and $\phi(y)=y_{1}.$ Let $c_{2}:=t/\phi(t).$ If $s^{\prime}=1$ then let $c_{1}:=s-\phi(s)$ and if $s^{\prime}=cs$ then let $c_{1}:=s/\phi(s).$ In any event, note that $c_{1},c_{2}$ are constants and that $s,t\in C(x_{1},y_{1},c_{1},c_{2}).$ Since $x,y$ are algebraic over $C(s,t),$ they are also algebraic over $C(x_{1},y_{1},c_{1},c_{2})$. This shows that the system (A) is almost internal to the constants. ∎
We conclude this article with the following observations. When $g(x)=x,$ we recover Theorem B of [JM20]. The nonorthogonality of the system (A), under the assumption that $x^{\prime}=f(x)$ is orthogonal to the constants, has been examined in [JJP23, Corollary 5.5]. Interestingly, the condition (ii) of Theorem 5 also appears in their result.
Acknowledgement. We thank the anonymous referee of this article for various helpful suggestions and references.